OpenStudy (anonymous):

Solve the equation using the Quadratic formula 11x^2=8

OpenStudy (kenljw):

11x^2 = 8 x^2 = 8/11 x = + sqrt(8/11) : - sqrt(8/11) x = + 2 sqrt(2/11) : - 2 sqrt(2/11)